# and even simpler proofs of quantumness

### v1

@@ 2.6 Contribution 2 — An even simpler proof of quantumness
Figure 6: The prover’s circuit. Here, $(f_{0},f_{1})$ is a pair of trapdoor claw-free functions (with inputs of size $n$). $U_{f_{0},f_{1}}$ denotes the $(n+1)$-qubit unitary that coherently computes $f_{0}$ in the last $n$ qubits if the first qubit is |$\ket{0}$, and computes $f_{1}$ otherwise. The circuit starts by preparing $\ket{+}$ in the first qubit, and a uniform superposition over all inputs in the next $n$ qubits. The circuit then applies $U_{f_{0},f_{1}}$ (note that we are omitting auxiliary work registers that are required to compute $U_{f_{0},f_{1}}$), followed by a layer of Hadamard gates on the last $n-1$ qubits. Then, the last $n$ qubits are measured. As a result, the leftover qubit |$\ket{\psi}$ in the first register is now a BB84 state (which one it is depends on $f_{0}$, $f_{1}$, and the measurement outcome). As its second message, the verifier sends a bit $b$, and the prover applies the rotation $R_{b}$ defined as follows: |$\ket{0}\overset{R_{b}}{\mapsto}\cos((-1)^{b}\pi/8)\ket{0}+\sin((-1)^{b}\pi/8)\ket{1}$ and |$\ket{1}\overset{R_{b}}{\mapsto}-\sin((-1)^{b}\pi/8)\ket{0}+\cos((-1)^{b}\pi/8)\ket{1}$. Finally, the prover measures the qubit in the standard basis. [^18] The simplified description of the proof of quantumness before this figure is slightly inaccurate: it states that the prover prepares starts by preparing the two-qubit state |+|$\ket{+}\ket{0}$. Technically, the prover only needs to prepare $\ket{+}$, and the role of the second qubit is performed by the first qubit of the pre-image register (which is initialised as a uniform superposition). We refer to Section 4 for a detailed description.
@@
An efficient quantum prover can efficiently pass this test with probability $\cos^{2}(\frac{\pi}{8})\approx 0.85$, while an efficient classical prover can pass this test with probability at most $3/4$. The proof of classical soundness is fairly straightforward. In essence, a classical prover can be rewound to obtain answers to both of the verifier’s possible questions in Message 3. If the classical prover passes the test with probability $3/4+\delta$ (which is on average over the two possible questions), then the answers to both questions together must reveal some information about the encrypted bit $a$ (this is a simple consequence of how the CHSH winning conditions is defined). In particular, such a classical prover can be used to guess $a$ with probability $\frac{1}{2}+2\delta$. This breaks the security of the encryption, as long as $\delta$ is non-negligible. We defer the reader to Section 5.2 for more details.
@@ 3.2 QFHE Scheme
For the rest of this work, we restrict to QFHE schemes where encryption takes the following form. For an $n$-qubit state |$\ket{\psi}$,
@@
where $\{U_{k}\}_{k\in K}$ forms a group (potentially up to global phases) and $K(n)$ is a finite set. Furthermore, $\hat{k}$ is a classical encryption of $k$, which can be decrypted using a PPT algorithm (specified by the $\mathsf{QFHE}$ scheme) given $\mathsf{sk}$.
All known $\mathsf{QFHE}$ schemes satisfying Definition 4 [Mah20, Bra18], also satisfy the property above with $U_{k}$ being a Pauli pad and $k=(x,z)\in\{0,1\}^{2n}$ encoding which Pauli operator is applied, i.e. $U_{k}=X^{x}Z^{z}$.
@@ 3.3 Noisy Trapdoor Claw-Free Functions
$\Pr[s\leftarrow\mathcal{A}(\mathsf{pk}):(\mathsf{pk},\mathsf{sk})\leftarrow{\cal F}.\mathsf{Gen}(1^{\lambda},s)\,\,,\,\,s\leftarrow\{0,1\}]\leq\mathsf{negl}(\lambda)\,.$ (7)
@@
Part I Even Simpler Proof of Quantumness
@@ 6 Contextuality Games
For the following three definitions, let $\mathsf{G}=(Q,A,C^{{\rm all}})$ be a contextuality game.
Definition 12 (Classical/Non-Contextual Strategy for $\mathsf{G}$${\rm Strat}_{NC}(\mathsf{G})$).
@@
where by $\mathbf{a_{joint}}[C]=\mathbf{a}$ we mean that $\mathbf{a_{joint}}[\tilde{q}]=\mathbf{a}[\tilde{q}]$ for all $\tilde{q}\in C$ [^19] Note that $\mathbf{a_{joint}}$ is a vector of length $|Q|$ while $\mathbf{a}$ is a vector of length $|C|$. The set of all non-contextual strategies is denoted by ${\rm Strat}_{NC}(\mathsf{G})$.
We now define a quantum strategy for $\mathsf{G}$. It would be helpful to first define a “con-instance” corresponding to a contextuality game $\mathsf{G}$.
Definition 13 (Con-instance of $\mathsf{G}$).
A con-instance is simply a triple $({\cal H},\left|\psi\right\rangle,\mathbf{O})$, where ${\cal H}$ is a Hilbert space, $\left|\psi\right\rangle\in{\cal H}$ is a quantum state and $\mathbf{O}$ is a list of Hermitian operators (observables), indexed by the questions $Q$ (i.e. $\mathbf{O}[q]$ is an observable for each $q\in Q$) acting on ${\cal H}$. We call the triple $({\cal H},\left|\psi\right\rangle,\mathbf{O})$ a con-instance of $\mathsf{G}$ if $\mathbf{O}$ satisfies the following:
@@
Let $\mathbf{C}:=\{\{\mathbf{O}[q]\}_{q\in C}\}_{C\in C^{{\rm all}}}$ denote the set of contexts, except this time, we consider the operators corresponding to the questions. We refer to the list $({\cal H},\left|\psi\right\rangle,\mathbf{O},\mathbf{C})$ as an (explicit) con-instance of $\mathsf{G}$.
To be concrete, consider the Magic square as an example of a con-instance.
Example 5 (A con-instance for the magic square).
An explicit con-instance corresponding to LABEL:Exa:MagicSquare for the magic square is given by (,|$(\cal{H},\ket{\psi},\mathbf{O},\mathbf{C})$ where ${\cal{H}}:=\mathbb{C}^{2}\times\mathbb{C}^{2}$, |$\ket{\psi}\in\cal{H}$ is any fixed state, say |$\ket{\psi}=\ket{00}$, and $\mathbf{O}$ is specified by the following (indexed by the questions $Q$ in some canonical way):
$\begin{array}[]{ccc}\mathbb{I}\otimes X&Z\otimes\mathbb{I}&Z\otimes X\\ X\otimes\mathbb{I}&\mathbb{I}\otimes Z&X\otimes Z\\ X\otimes X&Z\otimes Z&XZ\otimes XZ\end{array}$ (17)
and ${\mathbf{C}}$ (set of “contexts”, i.e. commuting operators) is a set containing rows and columns of these operators, i.e. $\{\{\mathbb{I}\otimes X,X\otimes\mathbb{I},X\otimes X\},\ \{Z\otimes\mathbb{I},\mathbb{I}\otimes Z,Z\otimes Z\},\ \dots\{X\otimes X,Z\otimes Z,XZ\otimes XZ\}\}\in{\mathbf{C}}$.
Definition 14 (Quantum Strategy for $\mathsf{G}$${\rm Strat}_{Qu}(\mathsf{G})$).
A strategy is a quantum strategy, i.e. $P\in{\rm Strat}_{Qu}(G)$, iff there is a con-instance $({\cal H},\left|\psi\right\rangle,\mathbf{O})$ of $\mathsf{G}$ satisfying
@@
${{\rm valNC}}:=\max_{P\in{\rm Strat}_{NC}(\mathsf{G})}{\rm val}(P)\quad\text{and similarly}\quad{{\rm valQu}}:=\max_{P\in{\rm Strat_{Qu}(\mathsf{G})}}{\rm val}(P).$

### v2

@@ 2.6 Contribution 2 — An Even Simpler Proof of Quantumness
Figure 10: The prover’s circuit. Here, $(f_{0},f_{1})$ is a pair of trapdoor claw-free functions (with inputs of size $n$). $U_{f_{0},f_{1}}$ denotes the $(n+1)$-qubit unitary that coherently computes $f_{0}$ in the last $n$ qubits if the first qubit is |$\ket{0}$, and computes $f_{1}$ otherwise. The circuit starts by preparing $\ket{+}$ in the first qubit, and a uniform superposition over all inputs in the next $n$ qubits. The circuit then applies $U_{f_{0},f_{1}}$ (note that we are omitting auxiliary work registers that are required to compute $U_{f_{0},f_{1}}$), followed by a layer of Hadamard gates on the last $n-1$ qubits. Then, the last $n$ qubits are measured. As a result, the leftover qubit |$\ket{\psi}$ in the first register is now a BB84 state (which one it is depends on $f_{0}$, $f_{1}$, and the measurement outcome). As its second message, the verifier sends a bit $b$, and the prover applies the rotation $R_{b}$ defined as follows: |$\ket{0}\overset{R_{b}}{\mapsto}\cos((-1)^{b}\pi/8)\ket{0}+\sin((-1)^{b}\pi/8)\ket{1}$ and |$\ket{1}\overset{R_{b}}{\mapsto}-\sin((-1)^{b}\pi/8)\ket{0}+\cos((-1)^{b}\pi/8)\ket{1}$. Finally, the prover measures the qubit in the standard basis. [^22] The simplified description of the proof of quantumness before this figure is slightly inaccurate: it states that the prover prepares starts by preparing the two-qubit state |+|$\ket{+}\ket{0}$. Technically, the prover only needs to prepare $\ket{+}$, and the role of the second qubit is performed by the first qubit of the pre-image register (which is initialised as a uniform superposition).
@@
[…]
@@ 3.2 QFHE Scheme
For the rest of this work, we restrict to (public key) QFHE schemes where encryption takes the following form. For an $n$-qubit state |$\ket{\psi}$,
@@
where $\{U_{k}\}_{k\in K}$ forms a group (potentially up to global phases) and $K(n)$ is a finite set. Furthermore, $\hat{k}$ is a classical fully homomorphic encryption of $k$, which can be decrypted using a PPT algorithm (specified by the $\mathsf{QFHE}$ scheme) given $\mathsf{sk}$.
We show that such $\mathsf{QFHE}$ schemes also satisfy the following property.
Classical evaluation of classical circuits.
There is a PPT procedure $\mathsf{cEval}$ that is the classical version of $\mathsf{Eval}$. More precisely, we require that the following properties hold:
  • For any classical ciphertext $c$ produced by running $\mathsf{Enc}$ on a classical message $m$, and any classical circuit $C$, it holds that $c^{\prime}\leftarrow\mathsf{cEval}(C,c)$ is such that $C(m)=\mathsf{Dec}(\mathsf{sk},c)$.
  • $\mathsf{cEval}$ and $\mathsf{Eval}$ produce identical distributions when run on identical classical inputs (i.e. ciphertexts and circuits).
Claim 6 (Classical Eval).
Consider any (public-key) $\mathsf{QFHE}$ scheme whose ciphertext is of the form in Equation 5. Then, there exists a PPT procedure $\mathsf{cEval}$, satisfying the classical evaluation of classical circuits property stated above.
All known $\mathsf{QFHE}$ schemes satisfying Definition 5 [Mah20, Bra18] are public key schemes that satisfy both properties listed above. In particular, the first property above (form of encryption) is satisfied with $U_{k}$ being a Pauli pad and $k=(x,z)\in\{0,1\}^{2n}$ encoding which Pauli operator is applied, i.e. $U_{k}=X^{x}Z^{z}$. We end by outlining how the second property (classical evaluation of classical circuits) is proved generically as stated in LABEL:Claim:classicalEval.
Proof sketch for LABEL:Claim:classicalEval.
We will show how the classical evaluation works by looking at homomorphic evaluation of the ”xor” (addition) and ”and” (multiplication) operations, that are universal for classical circuits. Consider two ciphertexts $\tilde{c}_{0}$ and $\tilde{c}_{1}$ corresponding to two messages $m_{0}$ and $m_{1}$, i.e. $\tilde{c}_{0}$ and $\tilde{c}_{1}$ are of the form $\tilde{c}_{0}=(m_{0}\oplus k_{0},c_{k_{0}})$ and $\tilde{c}_{1}=(m_{1}\oplus k_{1},c_{k_{1}})$ where $c_{k_{0}}\leftarrow\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(k_{0})$ and $c_{k_{1}}\leftarrow\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(k_{1})$. Then to homomorphically evaluate the addition, $\mathsf{cEval}$ will just add $\tilde{c}_{0}$ and $\tilde{c}_{1}$, namely compute $\tilde{c}_{m_{0}\oplus m_{1}}=(m_{0}\oplus k_{0}\oplus m_{1}\oplus k_{1},c_{k_{0}}+c_{k_{1}})=((m_{0}\oplus m_{1})\oplus(k_{0}\oplus k_{1}),c_{k_{0}\oplus k_{1}})$, where $c_{k_{0}\oplus k_{1}}$ is the $\mathsf{FHE}$ encryption of $k_{0}\oplus k_{1}$.
For the homomorphic multiplication operation, we will first compute: $(m_{0}\oplus k_{0})\cdot(m_{1}\oplus k_{1})=(m_{0}\cdot m_{1})\oplus(k_{0}\cdot m_{1}\oplus k_{1}\cdot m_{0}\oplus k_{0}\cdot k_{1})$. As a result, to complete the homomorphic evaluation, it suffices to show how to compute a $\mathsf{FHE}$ encryption of the pad $k_{0}\cdot m_{1}\oplus k_{1}\cdot m_{0}\oplus k_{0}\cdot k_{1}$. This can be done as follows. First, compute the $\mathsf{FHE}$ encryptions $c_{m_{0}\oplus k_{0}}=\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(m_{0}\oplus k_{0})$ and $c_{m_{1}\oplus k_{1}}=\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(m_{1}\oplus k_{1})$. Then given $c_{k_{0}}$ and $c_{m_{0}\oplus k_{0}}$, the algorithm will homomorphically compute the xor of the two messages (under encryption), resulting in obtaining $c_{m_{0}}=\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(m_{0})$ (and analogously $c_{m_{1}}=\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(m_{1})$). Next, given $c_{k_{1}}$ and $c_{m_{0}}$ one can homomorphically compute $c_{m_{0}\oplus k_{1}}=\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(m_{0}\oplus k_{1})$ and similarly, $c_{m_{1}\oplus k_{0}}=\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(m_{1}\oplus k_{0})$.
The procedure will then sample a uniformly random $r$ and compute $c_{r}=\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(r)$.
Finally, by homomorphically adding all the pieces together we get the final ciphertext of the multiplication: $\tilde{c}_{m_{0}\cdot m_{1}}:=((m_{0}\cdot m_{1})\oplus(k_{0}\cdot m_{1}\oplus k_{1}\cdot m_{0}\oplus k_{0}\cdot k_{1}\oplus r),\mathsf{FHE}.\mathsf{Enc}_{\mathsf{pk}}(k_{0}\cdot m_{1}\oplus k_{1}\cdot m_{0}\oplus k_{0}\cdot k_{1}\oplus r))$. It is clear to see that the ciphertext is in the right $\mathsf{QFHE}$ form, and moreover, due to the uniform pad $r$ we also have that the corresponding key is uniform, hence, ensuring that the output of $\mathsf{cEval}$ is identically distributed to the output of $\mathsf{Eval}$.
@@ 3.3 Noisy Trapdoor Claw-Free Functions
We also implicitly require that the classical version of efficient range superposition holds: uniform elements from the set ${\cal X}$ (and therefore ${\cal D}_{\cal Y}$ and ${\cal Y}$) can be efficiently sampled classically. This is the case for most NTCFs.
@@
$\Pr[s\leftarrow\mathcal{A}(\mathsf{pk}):(\mathsf{pk},\mathsf{sk})\leftarrow{\cal F}.\mathsf{Gen}(1^{\lambda},s)\,\,,\,\,s\leftarrow\{0,1\}]\leq 1/2+\mathsf{negl}(\lambda)\,.$ (7)
@@
Part I Even Simpler Proofs of Quantumness
@@ 6 Contextuality Games
For the following two definitions, let $\mathsf{G}=(Q,A,C^{{\rm all}})$ be a contextuality game.
Definition 14 (Classical/Non-Contextual Strategy for $\mathsf{G}$${\mathcal{S}}_{NC}(\mathsf{G})$).
@@
where by $\mathbf{a_{joint}}[C]=\mathbf{a}$ we mean that $\mathbf{a_{joint}}[\tilde{q}]=\mathbf{a}[\tilde{q}]$ for all $\tilde{q}\in C$ [^23] Note that $\mathbf{a_{joint}}$ is a vector of length $|Q|$ while $\mathbf{a}$ is a vector of length $|C|$. The set of all non-contextual strategies is denoted by ${\mathcal{S}}_{NC}(\mathsf{G})$.
We now define a quantum strategy for $\mathsf{G}$. It would be helpful to first define “qstrat”—a state and collection of observables—consistent with a contextuality game $\mathsf{G}$.
Definition 15 (Quantum Strategy for $\mathsf{G}$${\mathcal{S}}_{Qu}(\mathsf{G})$).
Given a contextuality game $\mathsf{G}$, consider a triple
$\mathsf{qstrat}:=({\cal H},\left|\psi\right\rangle,\mathbf{O}),$
where ${\cal H}$ is a Hilbert space, $\left|\psi\right\rangle\in{\cal H}$ is a quantum state and $\mathbf{O}$ is a list of Hermitian operators (observables), indexed by the questions $Q$ (i.e. $\mathbf{O}[q]$ is an observable for each $q\in Q$) acting on ${\cal H}$. Furthermore, suppose that $\mathbf{O}$ satisfies the following:
@@
A strategy is a quantum strategy, i.e. $P\in{\mathcal{S}}_{Qu}(G)$, iff there is such a triple $\mathsf{qstrat}:=({\cal H},\left|\psi\right\rangle,\mathbf{O})$ satisfying
@@
To be concrete, consider the Magic square as an example.
Example 5 ($\mathsf{qstrat}$ for the magic square).
For the magic square as in LABEL:Exa:MagicSquare, the optimal quantum strategy corresponds to using $\mathsf{qstrat}:=(\cal{H},\ket{\psi},\mathbf{O})$ where ${\cal{H}}:=\mathbb{C}^{2}\times\mathbb{C}^{2}$, |$\ket{\psi}\in\cal{H}$ is any fixed state, say |$\ket{\psi}=\ket{00}$, and $\mathbf{O}$ is specified by the following (indexed by the questions $Q$ in some canonical way):
$\begin{array}[]{ccc}\mathbb{I}\otimes X&Z\otimes\mathbb{I}&Z\otimes X\\ X\otimes\mathbb{I}&\mathbb{I}\otimes Z&X\otimes Z\\ X\otimes X&Z\otimes Z&XZ\otimes XZ.\end{array}$ (18)
@@
${{\rm valNC}}:=\max_{P\in{\mathcal{S}}_{NC}(\mathsf{G})}{\rm val}(P)\quad\text{and similarly}\quad{{\rm valQu}}:=\max_{P\in{\mathcal{S}_{Qu}(\mathsf{G})}}{\rm val}(P).$